# Multi-Scale Contrastive Learning for Video Temporal Grounding

### v1

@@ Introduction
Contrastive Learning. Notable improvements contributed by contrastive learning are made by contrastive loss applied to the final encoded outputs (Nguyen and Luu 2021; Nguyen et al. 2025; Hu, Cui, and Wang 2021; Nguyen et al. 2022, 2024b; Wang et al. 2021b; Nguyen et al. 2023a; Wu et al. 2024; Nguyen et al. 2024a). Wang et al. (2021b); Hu, Cui, and Wang (2021) employ a memory bank to maintain an extended set of positive and negative samples. Instead of utilizing outputs at the single final layer, contrastive learning with local and global representations across different layers has been widely studied (Hjelm et al. 2018; Zhang et al. 2020c; Bachman, Hjelm, and Buchwalter 2019; Chaitanya et al. 2020). Zhang et al. (2020c) maximize the mutual information between representations of different local windows of a sentence and the representation of the global sentence. In addition, Bachman, Hjelm, and Buchwalter (2019) propose an InfoNCE loss which optimizes over local and global features of two augmented views of an image, whereas Chaitanya et al. (2020) specifically concentrate on medical images.
@@
To fairly compare with previous works and satisfy the scalability of temporal grounding operation for long videos, we adopt video-centric sampling approach (Mu, Mo, and Li 2024). For Ego4D-NLQ, we use pre-trained 1) SlowFast video features (Feichtenhofer et al. 2019) with BERT textual features (Devlin et al. 2018), and 2) EgoVLP video and textual features (Lin et al. 2022). For testing, we report R@$\{1,5\}$, tIoU = $\{0.3,0.5\}$. For MAD dataset, we use CLIP features (Radford et al. 2021) for both videos and texts, and report R@$\{1,5,10,50\}$, tIoU = $\{0.1,0.3,0.5\}$. For the TACoS dataset, we use C3D video features (Tran et al. 2015) and GloVe textual features (Pennington, Socher, and Manning 2014). We report results in terms of R@$\{1,5\}$, tIoU = $\{0.5,0.7\}$. In addition, we utilize I3D features (Carreira and Zisserman 2017) pre-trained on Kinetics (Kay et al. 2017) for Charades-STA and C3D features (Tran et al. 2015) for ActivityNet-Captions experiments. For both datasets, similar to TACoS, we take advantage of GloVe textual features (Pennington, Socher, and Manning 2014). We report R@$\{1,5\}$, tIoU = $\{0.5,0.7\}$ for testing on Charades-STA, and R@$\{1,5\}$, tIoU = $\{0.3,0.5\}$ for testing on ActivityNet-Captions. For more details regarding model architecture, we direct interested readers to Appendix B. For both within-scale and cross-scale contrastive learning implementation, we keep the size of the negative sample set $\mathcal{N}(l)$ in every level $l$ to be equal to the size of the positive video clips $\mathcal{P}(l)$ that correspond to the target video moments. Based upon validation and fair comparison with previous methods, we use $\rho_{\text{ref}}=\rho_{\text{within}}=\rho_{\text{cross}}$ = 1.0.
@@
We consider the following temporal grounding models as baselines:
  • VSL-Net (Zhang et al. 2020a): utilizes textual query to highlight regions potential to comprise the target moment.
  • VLG-Net (Soldan et al. 2021): models temporal grounding as a graph matching problem.
  • Moment-DETR (Lei, Berg, and Bansal 2021): a Transformer encoder-decoder architecture that views temporal grounding as a set prediction problem.
  • CONE (Hou et al. 2022): subsequently slices a video input into windows, selects relevant windows, and ranks the selected windows to obtain target moments.
  • MMN (Wang et al. 2022): a Siamese-like network architecture that is trained with video-query and query-video contrastive learning.
  • SSRN (Zhu et al. 2023): enriches anchor frames with additional consecutive frames.
  • G2L (Li et al. 2023): measures moment-query similarities using geodesic distance and quantifies cross-modal interactions with game-theoretic interactions.
  • SOONet (Pan et al. 2023): an anchor-based framework that conducts grounding by pre-ranking, re-ranking, and regression.
  • MESM (Liu et al. 2024): a fine-grained moment-query contrastive approach modeled for query word and video moment representations,
  • Contrastive-MSAT (Panta et al. 2024): applies moment-query contrastive loss supported by a momentum-based memory bank.
  • UVCOM (Xiao et al. 2024): a moment-query contrastive approach for a unified video comprehension framework,
  • SnAG (Mu, Mo, and Li 2024): achieves scalable grounding with cross-modal late fusion.
@@ Acknowledgements
Appendix A More Qualitative Examples
We provide additional temporal grounding examples. As shown in Figure 4, our model can localize target video moments regardless of its lengths. In contrast, due to degraded information at higher levels of the feature pyramid, previous method ineffectively detects the moment given input queries.
Figure 4: More qualitative temporal grounding results.
Appendix B Model Architecture Details
Dataset Video Text Training
FPS Feature Embedding dim Window size $W$ Feature pyramid layers $L$ Feature Embedding dim Encoder layers Batch size Learning rate Epochs
MAD 5 CLIP 512 17 9 CLIP 512 1 32 1e-4 8
Ego4D-NLQ 30 SlowFast 512 19 8 BERT 128 1 2 1e-4 10
EgoVLP 384 EgoVLP 768 7
TACoS 29.4 C3D 128 19 8 GloVe 128 5 4 1e-3 15
Charades-STA 24 I3D 256 5 7 GloVe 128 5 4 1e-3 10
ActivityNet-Captions varying C3D 128 5 7 GloVe 128 5 2 1e-3 15
Table 7: Hyperparameter details of our model architecture.
In this appendix, we delineate our model architecture details. Table 7 summarizes key hyperparameters related to our architecture. In particular, the embedding dimension, the window size of the local self-attention, and the number of levels in video encoder’s feature pyramid differ across datasets. For the text encoder, we utilize a 1-layer Transformer encoder for MAD and Ego4D-NLQ datasets, and a 5-layer encoder for other datasets. All of the attention layers in our Transformer-based modules exhibit 4 attention heads. In our moment decoder, the 1D convolutional layers of the classification and regression heads possess a kernel of size 3.

### v2

@@ Introduction
Contrastive Learning. Notable improvements contributed by contrastive learning are made by contrastive loss applied to the final encoded outputs (Nguyen and Luu 2021; Wu et al. 2023; Nguyen et al. 2025; Hu, Cui, and Wang 2021; Nguyen et al. 2022; Wu et al. 2024; Nguyen et al. 2024b; Wang et al. 2021b; Nguyen et al. 2023a, 2024a). Wang et al. (2021b); Hu, Cui, and Wang (2021) employ a memory bank to maintain an extended set of positive and negative samples. Instead of utilizing outputs at the single final layer, contrastive learning with local and global representations across different layers has been widely studied (Hjelm et al. 2018; Zhang et al. 2020c; Bachman, Hjelm, and Buchwalter 2019; Chaitanya et al. 2020). Zhang et al. (2020c) maximize the mutual information between representations of different local windows of a sentence and the representation of the global sentence. In addition, Bachman, Hjelm, and Buchwalter (2019) propose an InfoNCE loss which optimizes over local and global features of two augmented views of an image, whereas Chaitanya et al. (2020) specifically concentrate on medical images.
@@
To fairly compare with previous works and satisfy the scalability of temporal grounding operation for long videos, we adopt video-centric sampling approach (Mu, Mo, and Li 2024). For Ego4D-NLQ, we use pre-trained 1) SlowFast video features (Feichtenhofer et al. 2019) with BERT textual features (Devlin et al. 2018), and 2) EgoVLP video and textual features (Lin et al. 2022). For testing, we report R@$\{1,5\}$, tIoU = $\{0.3,0.5\}$. For MAD dataset, we use CLIP features (Radford et al. 2021) for both videos and texts, and report R@$\{1,5,10,50\}$, tIoU = $\{0.1,0.3,0.5\}$. For the TACoS dataset, we use C3D video features (Tran et al. 2015) and GloVe textual features (Pennington, Socher, and Manning 2014). We report results in terms of R@$\{1,5\}$, tIoU = $\{0.5,0.7\}$. In addition, we utilize I3D features (Carreira and Zisserman 2017) pre-trained on Kinetics (Kay et al. 2017) for Charades-STA and C3D features (Tran et al. 2015) for ActivityNet-Captions experiments. For both datasets, similar to TACoS, we take advantage of GloVe textual features (Pennington, Socher, and Manning 2014). We report R@$\{1,5\}$, tIoU = $\{0.5,0.7\}$ for testing on Charades-STA, and R@$\{1,5\}$, tIoU = $\{0.3,0.5\}$ for testing on ActivityNet-Captions. For more details regarding model architecture, we direct interested readers to the appendix. For both within-scale and cross-scale contrastive learning implementation, we keep the size of the negative sample set $\mathcal{N}(l)$ in every level $l$ to be equal to the size of the positive video clips $\mathcal{P}(l)$ that correspond to the target video moments. Based upon validation and fair comparison with previous methods, we use $\rho_{\text{ref}}=\rho_{\text{within}}=\rho_{\text{cross}}$ = 1.0.
@@
We consider the following temporal grounding models as baselines: (i) VSL-Net (Zhang et al. 2020a) utilizing textual query to highlight regions potential to comprise the target moment; (ii) VLG-Net (Soldan et al. 2021) modeling temporal grounding as a graph matching problem; (iii) Moment-DETR (Lei, Berg, and Bansal 2021), a Transformer encoder-decoder architecture that views temporal grounding as a set prediction problem; (iv) CONE (Hou et al. 2022) subsequently slicing a video input into windows, selects relevant windows, and ranks the selected windows to obtain target moments; (v) MMN (Wang et al. 2022), a Siamese-like network architecture that is trained with video-query and query-video contrastive learning; (vi) SSRN (Zhu et al. 2023) enriching anchor frames with additional consecutive frames; (vii) G2L (Li et al. 2023) measuring moment-query similarities using geodesic distance and quantifies cross-modal interactions with game-theoretic interactions; (viii) SOONet (Pan et al. 2023), an anchor-based framework that conducts grounding by pre-ranking, re-ranking, and regression; (ix) MESM (Liu et al. 2024), a fine-grained moment-query contrastive approach modeled for query word and video moment representations; (x) Contrastive-MSAT (Panta et al. 2024), applying moment-query contrastive loss supported by a momentum-based memory bank; (xi) UVCOM (Xiao et al. 2024), a moment-query contrastive approach for a unified video comprehension framework; (xii) SnAG (Mu, Mo, and Li 2024) achieving scalable grounding with cross-modal late fusion.